OpenStudy (anonymous):

Simplify. Show your work. 5 1/3 +(-3 9/18)

OpenStudy (whpalmer4):

How would you write 1/3 as x/18? Figure that out, then you can add everything together

OpenStudy (anonymous):

multiply the denominator by 6?

hero (hero):

First simplify the fraction 9/18 to 1/2. Then write your fractional expression as: \[\space\space\space\space\space5\frac{1}{3} \\ - 3 \frac{1}{2} \\ \text{_______}\]

OpenStudy (anonymous):

Is this what your question look like? \[5\frac{ 1 }{ 3 } + (-3\frac{ 9 }{ 18 })\]

OpenStudy (anonymous):

@incomplte yes ^^

hero (hero):

@she-walks-in, did you read anything I wrote at all?

OpenStudy (anonymous):

@Hero yes I just saw it and I get it now, 9/18 simplifies to 1/2. Now how do i subtract 3 1/2 from 5 1/3. The denominators are still different. or is that just for multiplication that i have to have that?

OpenStudy (anonymous):

Ok. I would advise you to first multiply out into standard fraction. The first part would become 16/3. What do you think the second part that you already simplified would become?

hero (hero):

@incomplte, I got this, okay?

OpenStudy (anonymous):

Sorry i was just trying to help too, but of course go ahead.

OpenStudy (anonymous):

thanks for your help =) @incomplte

hero (hero):

Well, that's because you're explaining a different method from the one that I am showing her. When that happens, things can get confusing

hero (hero):

I started helping her before you came so be respectful and let me continue

hero (hero):

So as I was saying @she-walks-in, we have: \[\space\space\space\space\space5\frac{1}{3} \\ - 3 \frac{1}{2} \\ \text{_______}\]

hero (hero):

Now we have to find equivalent fractions for 1/3 and 1/2 such that they have the same denominator. To do that, we first must find the LCM between 2 and 3.

OpenStudy (anonymous):

The LCM would be 6, right?

hero (hero):

Precisely. Do you have an idea of what the equivalent fractions might be for 1/3 and 1/2?

hero (hero):

Basically, all you would have to do is find two numbers x and y: \[\frac{1}{3} = \frac{x}{6}\] \[\frac{1}{2} = \frac{y}{6}\]

hero (hero):

such that the equations above are equal.

OpenStudy (anonymous):

proportions?

hero (hero):

Yes, figure out what x and y are so that we can use them to finish doing the problem.

OpenStudy (anonymous):

x=2 y=3

hero (hero):

Okay, good. So now we have \(\frac{2}{6}\) and \(\frac{3}{6}\)

hero (hero):

So we will replace those in the original problem to get \[\space\space\space\space\space5\frac{2}{6} \\ - 3 \frac{3}{6} \\ \text{_______}\]

hero (hero):

Now before we subtract, we have something else we must do, which is to borrow one from 5 so that we can subtract the fractions appropriately.

hero (hero):

Borrowing one from 5 would make the fraction at the top look something like this: \[4\frac{2}{6} + \frac{6}{6} = 4\frac{8}{6}\]

hero (hero):

I hope that makes sense. If not, I can explain it.

hero (hero):

Ultimately, what we need to subtract here is: \[\space\space\space\space\space4\frac{8}{6} \\ - 3 \frac{3}{6} \\ \text{_______}\]

hero (hero):

@she-walks-in, are you stuck or have you figured it out yet?

OpenStudy (anonymous):

sorry my computer randomly blue screened and shut off. I'm confused how you got from the answers we go from the proportions to this

hero (hero):

I hope you at least understand how we got 2/6 and 3/y6

OpenStudy (anonymous):

ok, sorry i just had to reread it. I get it.

hero (hero):

So.....did you come up with a result? If so, I'd like to know.

OpenStudy (anonymous):

17/6 on my calculater

hero (hero):

On your calculator? Really? I'm sad that you didn't simply subtract the fraction we came up with.

hero (hero):

The calculator will give you the improper form

OpenStudy (anonymous):

erm 11/6 i mean

OpenStudy (anonymous):

and yup so that'll be 1 5/6 ?

hero (hero):

Yes, but you're supposed to use your calculator only for verification, and AFTER you have solved the problem. I'm kinda disappointed in you.

hero (hero):

\[\space\space\space\space\space4\frac{8}{6} \\ - 3 \frac{3}{6} \\ \text{_______} \\ \space\space \space \space 1\frac{5}{6}\]

OpenStudy (anonymous):

sorry =/ My teachers never really told me that..um.. I'll try not to do that so much anymore

hero (hero):

You must be one of those online homeschoolers

OpenStudy (anonymous):

yessir, only just. I've just started this september

OpenStudy (anonymous):

well, thank you anyway, i really do appreciate it